# references.bib

### v2

@@ 2.2. Group actions on handlebodies
Using the Equivariant Loop Theorem [meeks-yau-loop], we can parition $V$ into $G$-orbits of 00-handles and $1$-handles. For each $g\in G$ and 00-handle $E\cong B^{3}$, either $g(E)\cap E=\varnothing$ or $g$ acts on $E$ by an element of $\operatorname{SO}(3)$ (any finite group action on the $3$-ball is equivalent to a linear action [kwasik-schultz-actions-on-3ball]). For each $g\in G$ and $1$-handle $F\cong D^{2}\times I$, either $g(F)\cap F=\varnothing$ or $g$ rotates the $D^{2}$-factor of $F$. Note that since $V$ was constructed as a neighborhood of $B$, every 00-handle and $1$-handle has points fixed by some $g\in G$.
@@ 4.1. Slide homeomorphisms
We define the notion of a slide homeomorphism following McCullough [mccullough, §1]. Recall that the base point $*\in N^{\circ}$ lies on $S$. Let $\gamma$ be a simple closed curve in $N_{i}^{\circ}$ based at *. View $W^{\circ}_{i}$ as obtained from capping the boundary sphere $S$ of $N_{i}^{\circ}$ by a 3-ball $E$. Let $h_{t}$ be an ambient isotopy of $W^{\circ}_{i}$ with $h_{0}=\operatorname{id}$ and $h_{1}|_{E}=\operatorname{id}_{E}$ which moves $E$ along $\gamma$. Then, we define a slide along $\gamma$ to be a homeomorphism which acts on $N_{i}^{\circ}$ by $h_{1}$ for some such isotopy $h_{t}$, and which fixes the other summand $N_{j}^{\circ}$ pointwise. We warn the reader that for two choices of isotopy $h_{t}$ and $h_{t}^{\prime}$, the resulting slide homeomorphisms may differ up to isotopy by a sphere twist about $S$.
@@
Suppose that $\gamma$ is a simple closed curve in $N_{1}^{\circ}$ based at * and $f$ is a slide along $\gamma$. Since $f$ fixes *, it induces an automorphism of $\pi_{1}(N^{\circ})$. In particular, we see that $f$ acts on $\pi_{1}(N^{\circ})$ by partial conjugation. That is, with respect to the free splitting $\pi_{1}(N^{\circ})=\pi_{1}(N_{1}^{\circ})*\pi_{1}(N_{2}^{\circ})$, the homeomorphism $f$ conjugates $\pi_{1}(N_{1}^{\circ})$ by $[\gamma]$ and fixes $\pi_{1}(N_{2}^{\circ})$ pointwise. The analogous statement holds if $\gamma$ is a simple closed curve in $N_{2}^{\circ}$. This observation has the following immediate consequence.
@@ 4.2. Part (i) of Theorem A
Now, we prove part (i) of A by using Lemmas 4.1 and 4.2 to find an infinite order element of $\operatorname{Ker}(\mathcal{L}_{p})$.
@@ 4.3. Part (ii) of Theorem A
Now, combining Lemmas 4.4 and 4.5, we can prove part (ii) of Theorem A.
@@
First, suppose that $N_{i}^{\circ}$ has a curve $\gamma$ based at * which is freely homotopic to a branch meridian, and that the element $[\gamma]$ has no central power in $\pi_{1}(N_{i}^{\circ})$. Then, as in the proof of Proposition 4.3, we deduce from Lemmas 4.1 and 4.2 that $\operatorname{Ker}(\mathcal{L}_{p})$ is infinite.
@@ 6.2. The restriction map
Next, suppose that $\sigma\in S_{n}$. Then we can write $\sigma=\sigma^{\prime}$ or $\sigma=(i\ n)\circ\sigma^{\prime}$ for some $\sigma^{\prime}\in S_{n-1}$ and $i<n$. The element $\sigma^{\prime}$ acts on $F_{n-1}$ by permuting the indices of the generators $x_{1},\ldots,x_{n-1}$. The element $(i\ n)$ maps $x_{i}$ to $x_{i}^{-1}$ and each other $x_{j}$ to $x_{j}x_{i}^{-1}$. Thus, if $\sigma$ is nontrivial, then $s(\sigma)$ is nontrivial.
@@ 7.1. The simplicial complexes
We begin by defining the simplicial complexes, following the work of McCullough-Miller [mccullough-miller]. We note that McCullough-Miller primarily work with a definition in terms of “based partitions”; however, we will use their alternative definition in terms of “labelled bipartite trees” [mccullough-miller, §1.3], as this definition is similar to more familiar complexes like the Culler-Vogtman Outer space. We also note that McCullough-Miller discuss arbitrary free products $G_{1}*\cdots*G_{n}$, where $G_{i}$ is any freely indecomposable group; we are only concerned with the case that $G_{i}=\mathbb{Z}$ or $G_{i}=\mathbb{Z}/2\mathbb{Z}$, which will slightly simplify the dicussion.
@@ 7.2. Relating the complexes
Finally, we can prove Proposition 7.1. The main result of [mccullough-miller] is that the complexes $K_{0}(F_{n})$ and $K_{0}(H_{n})$ are contractible. By Lemma 7.3, $\operatorname{Ker}(\mathcal{P}_{n})$ acts on the contractible complex $K_{0}(F_{n})$ and the quotient is contractible. Moreover, we claim that $\operatorname{Ker}(\mathcal{P}_{n})$ acts without rotations, meaning that if $\alpha\in\operatorname{Ker}(\mathcal{P}_{n})$ preserves a simplex $s$ of $K_{0}(F_{n})$, then $\alpha$ fixes each vertex of $s$. Indeed, as we remarked in Section 7.1, any simplex contains at most one labelled bipartite tree of any given type; since $\operatorname{SymOut}(F_{n})$ preserves the type of a labelled bipartite tree, this implies that $\operatorname{SymOut}(F_{n})$ acts without rotations, and hence $\operatorname{Ker}(\mathcal{P}_{n})$ does too. Thus, a result of Armstrong [armstrong] says that $\operatorname{Ker}(\mathcal{P}_{n})$ is generated its stabilizers of vertices of $K_{0}(F_{n})$ (see also [putman-choices]). This implies $\operatorname{Ker}(\mathcal{P}_{n})$ is normally generated over $\operatorname{SymOut}(F_{n})$ by the $\operatorname{Ker}(\mathcal{P}_{n})$-stabilizers of one vertex from each $\operatorname{SymOut}(F_{n})$-orbit.
@@ 8.2. The case of two factors
From Lemma 8.1, we see that $\rho_{1}$ and $\rho_{2}$ both lift to $T$. We claim that $\sigma$ lifts to a homeomorphism which is isotopic to $\tau$. Indeed, let $\theta\in\operatorname{Homeo}^{+}(S^{1}\times S^{2})$ be a the $\pi$ rotation in the $S^{1}$ factor, and let $\tau^{\prime}=\theta\circ\tau$. Then $\tau^{\prime}$ is symmetric since $\theta$ commutes with $\tau$, and it descends to a homeomorphism of $(S^{3},C_{2})$ which permutes the components of $C_{2}$, which must represent $\sigma$.

### v3

@@ 1.5. Proof idea of Theorem A
Slide homeomorphisms can also be defined for surfaces, but they behave quite differently in the $3$-manifold case compared to the surface case. This difference is rooted in the fact that sphere twists on $3$-manifolds behave much differently from Dehn twists on surfaces (namely, sphere twists have order $2$ and act trivially on $\pi_{1}$). This difference helps explain why slide homeomorphisms provide elements in the kernel of the lifting map for $3$-manifolds, but not for surfaces. See Remark 4.2 for more details.
@@ 2.2. Group actions on handlebodies
Using the Equivariant Loop Theorem [meeks-yau-loop], we can partition $V$ into $G$-orbits of 00-handles and $1$-handles. For each $g\in G$ and 00-handle $E\cong B^{3}$, either $g(E)\cap E=\varnothing$ or $g$ acts on $E$ by an element of $\operatorname{SO}(3)$ (any finite group action on the $3$-ball is equivalent to a linear action [kwasik-schultz-actions-on-3ball]). For each $g\in G$ and $1$-handle $F\cong D^{2}\times I$, either $g(F)\cap F=\varnothing$ or $g$ rotates the $D^{2}$-factor of $F$. Note that since $V$ was constructed as a neighborhood of $B$, every 00-handle and $1$-handle has points fixed by some $g\in G$.
@@ 4.1. Slide homeomorphisms
We define the notion of a slide homeomorphism following McCullough [mccullough, §1], though we will take a slightly more formal approach. Recall that the base point $*\in N^{\circ}$ lies on $S$. Let $\gamma$ be a simple closed curve in $N_{i}^{\circ}$ based at *. View $W^{\circ}_{i}$ as obtained from capping the boundary sphere $S$ of $N_{i}^{\circ}$ by a 3-ball $E$.
First, we introduce the notion of a boundary slide of $N_{i}^{\circ}$ along $\gamma$. Informally, let $h_{t}$ be an ambient isotopy of $W_{i}^{\circ}$ with $h_{0}=\operatorname{id}$ and $h_{1}|_{E}=\operatorname{id}_{E}$ which moves $E$ along $\gamma$; a boundary slide is the restriction of $h_{1}$ to $N_{i}^{\circ}$ for any choice of such an isotopy $h_{t}$. We formalize this idea as follows (for convenience, we work in the smooth category). Choose an arc in $E$ from the center of $E$ to *, and conjugate $\gamma$ by this arc to obtain a loop $\gamma^{\prime}$ based at the center of $E$. Then choose a lift of $\gamma^{\prime}$ to a loop $\widetilde{\gamma}^{\prime}$ in the (oriented) frame bundle $\operatorname{Fr}^{+}(W_{i}^{\circ})$. The frame bundle $\operatorname{Fr}^{+}(W_{i}^{\circ})$ is weakly homotopy equivalent to the space $\operatorname{Emb}^{+}(E,W_{i}^{\circ})$ of smooth orientation-preserving embeddings of $E$ in $W_{i}^{\circ}$ [ivanov, Thm 2.6.C]. Moreover, there is a fibration
$\operatorname{Diff}_{\partial}^{+}(W_{i}^{\circ}\operatorname{rel}E)\rightarrow\operatorname{Diff}_{\partial}^{+}(W_{i}^{\circ})\rightarrow\operatorname{Emb}^{+}(E,W_{i}^{\circ}),$
where $\operatorname{Diff}_{\partial}^{+}(W_{i}^{\circ})$ is the group of orientation-preserving diffeomorphisms fixing $\partial W_{i}^{\circ}$ pointwise, and $\operatorname{Diff}_{\partial}^{+}(W_{i}^{\circ}\operatorname{rel}E)$ is the group of orientation-preserving diffeomorphisms that fix $\partial W_{i}^{\circ}$ and $E$ pointwise [ivanov, Thm 2.6.A]. We define a boundary slide of $N_{i}^{\circ}$ along $\gamma$ to be a homeomorphism which represents the image of $\widetilde{\gamma}^{\prime}$ under the composition
$\pi_{1}(\operatorname{Fr}^{+}(W_{i}^{\circ}))\xrightarrow{\cong}\pi_{1}(\operatorname{Emb}^{+}(E,W_{i}^{\circ}))\rightarrow\pi_{0}(\operatorname{Diff}_{\partial}^{+}(W_{i}^{\circ}\operatorname{rel}E))\rightarrow\pi_{0}(\operatorname{Diff}_{\partial}^{+}(N_{i}^{\circ})),$
where the second map comes from the long exact sequence on homotopy groups and the third map is induced by restriction.
The isotopy class (rel boundary) of a boundary slide depends only on the based homotopy class of the chosen lift $\widetilde{\gamma}^{\prime}$. Note that since any compact oriented $3$-manifold is parallelizable, the frame bundle is a trivial $\operatorname{GL}_{3}^{+}(\mathbb{R})$-bundle, and hence
$\pi_{1}(\operatorname{Fr}^{+}(W_{i}^{\circ}))\cong\pi_{1}(\operatorname{GL}_{3}^{+}(\mathbb{R}))\times\pi_{1}(W_{i}^{\circ})\cong\mathbb{Z}/2\mathbb{Z}\times\pi_{1}(W_{i}^{\circ}).$
Boundary slides coming the $\pi_{1}(W_{i}^{\circ})$-factor behave like “point pushing” maps on surfaces (see [farb-margalit, §4.2]), while a boundary slide coming from the $\pi_{1}(\operatorname{GL}_{3}^{+}(\mathbb{R}))$-factor is a sphere twist about $S$ (cf. [hatcher-wahl, Rem 2.4]). The based homotopy class of our lift $\widetilde{\gamma}^{\prime}$ is determined by the based homotopy class of $\gamma$ up to the $\pi_{1}(\operatorname{GL}_{3}^{+}(\mathbb{R}))$-factor. It follows that given an element of $\pi_{1}(N_{i}^{\circ},*)$, the isotopy class of the resulting boundary slide is well-defined up to a sphere twist about $S$.
Now, we define a slide along $\gamma$ to be a homeomorphism of $N^{\circ}$ which acts on $N_{i}^{\circ}$ by a boundary slide along $\gamma$, and which fixes the other summand $N_{j}^{\circ}$ pointwise. We again emphasize that given an element of $\pi_{1}(N_{i}^{\circ},*)$, the isotopy class of a resulting slide homeomorphism is only well-defined up to a sphere twist about $S$.
@@
Next, we can examine the action of a slide homeomorphism on $\pi_{1}(N^{\circ})$. Suppose that $\gamma$ is a simple closed curve in $N_{1}^{\circ}$ based at * and $f$ is a slide along $\gamma$. Since $f$ fixes *, it induces an automorphism of $\pi_{1}(N^{\circ})$. In particular, we see that $f$ acts on $\pi_{1}(N^{\circ})$ by partial conjugation. That is, with respect to the free splitting $\pi_{1}(N^{\circ})=\pi_{1}(N_{1}^{\circ})*\pi_{1}(N_{2}^{\circ})$, the homeomorphism $f$ conjugates $\pi_{1}(N_{1}^{\circ})$ by $[\gamma]$ and fixes $\pi_{1}(N_{2}^{\circ})$ pointwise. The analogous statement holds if $\gamma$ is a simple closed curve in $N_{2}^{\circ}$. This observation has the following immediate consequence.
@@
Remark 4.2.
Slide homeomorphisms have an analogous construction on surfaces. Let $\Sigma$ be a compact orientable surface, and let $\delta$ be a separating simple closed curve on $\Sigma$. Let $\Sigma_{1}$ and $\Sigma_{2}$ denote the closures of the components of $\Sigma-\delta$, and assume each $\pi_{1}(\Sigma_{i})$ is nontrivial. Let $\widehat{\Sigma}_{i}$ denote the surface obtained from $\Sigma_{i}$ by capping the boundary component coming from $\delta$ with a disk. We define a boundary slide of $\Sigma_{i}$ to be a homeomorphism of $\Sigma_{i}$ which represents an element of the kernel of the map $\operatorname{Mod}(\Sigma_{i})\rightarrow\operatorname{Mod}(\widehat{\Sigma}_{i})$. Boundary slides on $\Sigma_{i}$ correspond to elements of the fundamental group of the unit tangent bundle $\operatorname{UT}(\widehat{\Sigma}_{i})$ [farb-margalit, §4.2.5], just as boundary slides on $N_{i}^{\circ}$ correspond to elements of $\pi_{1}(\operatorname{Fr}^{+}(W_{i}^{\circ}))$. Then, we can define a slide homeomorphism of $\Sigma$ to be a homeomorphism which acts on $\Sigma_{i}$ by a boundary slide and which fixes the other subsurface $\Sigma_{j}$ pointwise.
For a surface $\Sigma$, the action of a slide homeomorphism on $\pi_{1}(\Sigma)$ is more complicated than in the $3$-manifold case. This is because $\pi_{1}(\operatorname{UT}(\widehat{\Sigma}_{i}))$ fits into a short exact sequence
$1\rightarrow\mathbb{Z}\rightarrow\pi_{1}(\operatorname{UT}(\widehat{\Sigma}_{i}))\rightarrow\pi_{1}(\widehat{\Sigma}_{i})\rightarrow 1$
Given a generator of the kernel $\mathbb{Z}$, the corresponding slide homeomorphism is a Dehn twist around $\delta$, which will act nontrivially on $\pi_{1}(\Sigma)$, unlike sphere twists in the $3$-manifold case. This short exact sequence also implies that an element of $\pi_{1}(\Sigma_{i})$ corresponds to infinitely many isotopy classes of slide homeomorphisms, unlike in the $3$-manifold case where there are at most $2$ isotopy classes. This second difference is ultimately a consequence of the fact that $\pi_{1}(\operatorname{GL}_{2}(\mathbb{R}))\cong\mathbb{Z}$ while $\pi_{1}(\operatorname{GL}_{3}(\mathbb{R}))\cong\mathbb{Z}/2\mathbb{Z}$.
@@ 4.2. Part (i) of Theorem A
Now, we prove part (i) of A by using Lemmas 4.1 and 4.3 to find an infinite order element of $\operatorname{Ker}(\mathcal{L}_{p})$.
@@ 4.3. Part (ii) of Theorem A
Now, combining Lemmas 4.5 and 4.6, we can prove part (ii) of Theorem A.
@@
First, suppose that $N_{i}^{\circ}$ has a curve $\gamma$ based at * which is freely homotopic to a branch meridian, and that the element $[\gamma]$ has no central power in $\pi_{1}(N_{i}^{\circ})$. Then, as in the proof of Proposition 4.4, we deduce from Lemmas 4.1 and 4.3 that $\operatorname{Ker}(\mathcal{L}_{p})$ is infinite.
@@ 6.2. The restriction map
Next, suppose that $\sigma\in S_{n}$. Note that if $\sigma(z_{n})=z_{n}$, then $\sigma$ simply permutes the indices of the generators $x_{1},\ldots,x_{n-1}$. Otherwise, we can write $\sigma=(i\ n)\circ\sigma^{\prime}$ for some $1\leq i<n$, where $(i\ n)$ is the transposition swapping $z_{i}$ and $z_{n}$, and $\sigma^{\prime}$ is some permutation that fixes $z_{n}$. The element $(i\ n)$ maps $x_{i}$ to $x_{i}^{-1}$ and each other $x_{j}$ to $x_{j}x_{i}^{-1}$, and $\sigma^{\prime}$ permutes the indices of the generators $x_{1},\ldots,x_{n-1}$. In any case, we see that if $\sigma$ is nontrivial, then $s(\sigma)$ is nontrivial.
@@ 7.1. The simplicial complexes
We begin by defining the simplicial complexes, following the work of McCullough-Miller [mccullough-miller]. We note that McCullough-Miller primarily work with a definition in terms of “based partitions”; however, we will use their alternative definition in terms of “labelled bipartite trees” [mccullough-miller, §1.3], as this definition is similar to more familiar complexes like the Culler-Vogtman Outer space. We also note that McCullough-Miller discuss arbitrary free products $G_{1}*\cdots*G_{n}$, where $G_{i}$ is any freely indecomposable group; we are only concerned with the case that $G_{i}=\mathbb{Z}$ or $G_{i}=\mathbb{Z}/2\mathbb{Z}$, which will slightly simplify the discussion.
@@ 7.2. Relating the complexes
Finally, we can prove Proposition 7.1. The main result of [mccullough-miller] is that the complexes $K_{0}(F_{n})$ and $K_{0}(H_{n})$ are contractible. By Lemma 7.3, $\operatorname{Ker}(\mathcal{P}_{n})$ acts on the contractible complex $K_{0}(F_{n})$ and the quotient is contractible. Moreover, we claim that $\operatorname{Ker}(\mathcal{P}_{n})$ acts without rotations, meaning that if $\alpha\in\operatorname{Ker}(\mathcal{P}_{n})$ preserves a simplex $s$ of $K_{0}(F_{n})$, then $\alpha$ fixes each vertex of $s$. Indeed, as we remarked in Section 7.1, any simplex contains at most one labelled bipartite tree of any given type; since $\operatorname{SymOut}(F_{n})$ preserves the type of a labelled bipartite tree, this implies that $\operatorname{SymOut}(F_{n})$ acts without rotations, and hence $\operatorname{Ker}(\mathcal{P}_{n})$ does too. Thus, a result of Armstrong [armstrong] says that $\operatorname{Ker}(\mathcal{P}_{n})$ is generated by its stabilizers of vertices of $K_{0}(F_{n})$ (see also [putman-choices]). This implies $\operatorname{Ker}(\mathcal{P}_{n})$ is normally generated over $\operatorname{SymOut}(F_{n})$ by the $\operatorname{Ker}(\mathcal{P}_{n})$-stabilizers of one vertex from each $\operatorname{SymOut}(F_{n})$-orbit.
@@ 8.2. The case of two factors
From Lemma 8.1, we see that $\rho_{1}$ and $\rho_{2}$ both lift to $T$. We claim that $\sigma$ lifts to a homeomorphism which is isotopic to $\tau$. Indeed, let $\theta\in\operatorname{Homeo}^{+}(S^{1}\times S^{2})$ be a $\pi$ rotation in the $S^{1}$ factor, and let $\tau^{\prime}=\theta\circ\tau$. Then $\tau^{\prime}$ is symmetric since $\theta$ commutes with $\tau$, and it descends to a homeomorphism of $(S^{3},C_{2})$ which permutes the components of $C_{2}$, which must represent $\sigma$.